OpenStudy (anonymous):

Which rectangular equation is equivalent to the polar equation r*sin(theta)=-2? A. y= 2x^2 B. y=-2 C. x=-2 D. x^2+y^2 = -2

OpenStudy (anonymous):

In Polar coordinates \[ x= r \cos(\theta)\\ y= r \sin(\theta) \]

OpenStudy (anonymous):

Got it. Thanks

OpenStudy (anonymous):

So what is the answer?

OpenStudy (anonymous):

YW

OpenStudy (anonymous):